# Joint Transmit and Pinching Beamforming for Pinching Antenna Systems (PASS): Optimization-Based or Learning-Based?

### v2

@@ IV-C2 Decoder Design
The decoder dynamically attends to relevant information in the final hidden state $\mathbf{Z}_{\mathrm{enc}}$ obtained by the encoder, and then predicts primal/dual variables based on a cross-attention mechanism. Let $\mathbf{Z}_{\mathrm{dec}}$ denote the prediction of the decoder. At each decoding layer, each attention head $i$ of the decoder generates the key matrix for $\mathbf{Z}_{\mathrm{dec}}$, while computing queue matrix and value matrix using $\mathbf{Z}_{\mathrm{enc}}$:
$\widetilde{\mathbf{Y}}_{i}^{K}\!=\!\mathbf{Z}_{\mathrm{dec}}\!\widetilde{\mathbf{W}}_{i}^{K},\quad\widetilde{\mathbf{Y}}_{i}^{q}\!=\!\mathbf{Z}_{\mathrm{enc}}\!\widetilde{\mathbf{W}}_{i}^{q},\quad\widetilde{\mathbf{Y}}_{i}^{V}\!=\!\mathbf{Z}_{\mathrm{enc}}\!\widetilde{\mathbf{W}}_{i}^{V}.$ (28)
Hence, the cross attention score between $\mathbf{Z}_{\mathrm{enc}}$ and $\mathbf{Z}_{\mathrm{dec}}$ is obtained by:
$\mathrm{head}_{i}\!\!=\!\!CrAttn\!\!\left(\widetilde{\mathbf{Y}}_{i}^{K},\!\widetilde{\mathbf{Y}}_{i}^{q},\!\widetilde{\mathbf{Y}}_{i}^{V}\right)\!=\!\text{softmax}\!\!\left(\!\frac{\widetilde{\mathbf{Y}}_{i}^{q}\!\!\left(\widetilde{\mathbf{Y}}_{i}^{K}\right)^{\top}}{\sqrt{N_{\mathrm{model}}}}\!\right)\!\widetilde{\mathbf{Y}}_{i}^{V}\!.$ (29)
@@
$\mathbf{Z}_{\mathrm{dec}}\leftarrow\mathbf{Z}_{\mathrm{dec}}+\text{MultiHead}\left(\widetilde{\mathbf{Y}}_{i}^{K},\widetilde{\mathbf{Y}}_{i}^{q},\widetilde{\mathbf{Y}}_{i}^{V}\right).$ (30)
@@
Parameters: Model dimension $N_{\text{model}}$, number of heads $I$, numbers of encoder and decoder layers $J_{\mathrm{E}}$ and $J_{\mathrm{D}}$
1:  Compute feature embedding $\mathbf{Z}_{\mathrm{enc}}$ by (23).
@@
3:     Compute queue $\mathbf{Y}_{i}^{Q}$, key $\mathbf{Y}_{i}^{K}$, and value $\mathbf{Y}_{i}^{V}$ for each attention head $i$ by (24).
4:     Obtain self-attention by (25) and update $\mathbf{Z}_{\mathrm{enc}}$ by (26).
@@
6:  Initialize decoder output by $\mathbf{Z}_{\mathrm{dec}}\leftarrow\mathbf{Z}_{0,\mathrm{out}}+\text{PE}$, where positional encoding $\text{PE}_{(pos,2i)}=\sin\left(\frac{pos}{10000^{2i/d_{\text{model}}}}\right)$, $\text{PE}_{(pos,2i+1)}=\cos\left(\frac{pos}{10000^{2i/d_{\text{model}}}}\right)$.
@@
9:     Update $\mathbf{Z}_{\mathrm{dec}}$ by (30).
@@
11:  Obtain PA deployment $\mathbf{x}_{\mathrm{end}}\in\mathbb{R}^{N\times 1}$, $\bm{\omega}\in\mathbb{R}^{N\times L}$, dual variable $\bm{\lambda}\in\mathbb{R}^{K\times 1}$, and power allocation $\bm{\mu}\in\mathbb{R}^{K\times 1}$ from $\mathbf{Z}_{\mathrm{dec}}=\big{[}\mathbf{x}_{\mathrm{end}},\bm{\omega},\bm{\lambda},\bm{\mu}\big{]}\!\in\mathbb{R}^{(N+M+2K)\times 1}$.
12:  Project $\mathbf{x}_{\mathrm{end}}$, $\bm{\omega}$ by (31), (33) and obtain $\mathbf{X}$.
@@ IV-C3 Constraint Guarantees
We present projection operations in this part to strictly ensure the constraints in (C1) - (C3). First, to satisfy the pinching range limitation (C2), we newly introduce an $N$-dimension vector $\mathbf{x}^{\mathrm{end}}=\big{[}x_{1}^{\mathrm{end}},x_{2}^{\mathrm{end}},\dots,x_{N}^{\mathrm{end}}\big{]}$ that assigns the positions of the last PA along each waveguide. Therefore, we can simply project $\mathbf{x}^{\mathrm{end}}$ into feasible region $L\Delta_{\min}\leqslant\mathbf{x}^{\mathrm{end}}\leqslant S_{\mathrm{PA}}$ using
$\mathbf{x}^{\mathrm{end}}\leftarrow L\Delta_{\min}+\mathrm{Sigmoid}\left(\mathbf{x}^{\mathrm{end}}\right)\left(S_{\mathrm{PA}}-L\Delta_{\min}\right),$ (31)
where $\mathrm{Sigmoid}(\cdot)$ is the Sigmoid activation function. To simultaneously guarantee the minimum spacing constraint (C1), we further predict the spacing $\bm{\omega}_{n}^{T}=\left[\omega_{n,1},\omega_{n,2},\dots,\omega_{n,L}\right]\in\mathbb{R}^{1\times L}$ between any two PAs along the $n$-th waveguide. Specifically, $\omega_{n,1}$ denotes the distance from the first PA $l=1$ to the feed point of each waveguide $n$, and $\omega_{n,l}=x_{n,l}-x_{n,l-1}$, $\forall l\geqslant 1$, indicates the mutual intervals between the remaining PAs. We introduce the following operation to guarantee the minimum spacing constraint (C1). Given a threshold $0<\epsilon\leqslant 1/L$ and a positive vector $\mathbf{z}=\left[z_{1},z_{2},\dots,z_{L}\right]$, $0\leqslant z_{l}\leqslant 1$, the projection operation
$\mathbf{z}_{\mathrm{proj}}=\mathcal{P}\left(\mathbf{z},\epsilon\right)=\epsilon+\frac{1-L\epsilon}{\sum_{l=1}^{L}z_{l}}\mathbf{z}$ (32)
ensures that $\mathbf{z}_{\mathrm{proj}}\geqslant\epsilon$ and $\sum\limits_{l=1}z_{\mathrm{proj},l}=1$. Therefore, we can ensure the minimum spacing constraints using
$\bm{\omega}_{n}\leftarrow x_{n}^{\mathrm{end}}~{}\mathcal{P}\left(\bm{\omega}_{n},\frac{\Delta_{\min}}{x_{n}^{\mathrm{end}}}\right),~{}\forall n\in\mathcal{N}.$ (33)
As a result, the position $x_{n,l}$ of the $l$-th PA at waveguide $n$ can be obtained by accumulating the intervals of previous $l$ PAs using $x_{n,l}=\sum_{i=1}^{l}\omega_{n,i},~{}\forall l\in\mathcal{L},~{}\forall n\in\mathcal{N}$.
@@
Algorithm 2 summarizes the computation procedure for the developed KDL-Transformer algorithm for predicting solutions. Based on the proposed structure, the encoder and the decoder will be jointly trained through end-to-end learning to minimize the loss function, i.e., the negative system sum rate.
@@
Performance Method MM-PDD KDL-Transformer KDL-Transformer-OCA KDL-ResNet Black-box Transformer Black-box ResNet Conventional MIMO
@@ V Simulation Result
In this section, we provide numerical results to demonstrate the effectiveness of the proposed PASS design and algorithms, where a BS serves $K=\left\{4,6\right\}$ single-antenna users in downlink MIMO communications. The system operates at $f=30$ GHz frequency band, and the wavelength is $\lambda_{f}=0.01$ meters. Users are randomly distributed in a range of $S_{\mathrm{x}}\times S_{\mathrm{y}}=20\times 10~{}\text{m}^{2}$ area. Both users and PASS have fixed heights, where $z_{k}^{\mathrm{U}}=0$, $\forall k\in\mathcal{K}$, and $h_{\mathrm{PA}}=2.5$ meter, $\forall n\in\mathcal{N}$. To ensure multiplexing gains, the number of RF chains used by PASS is set as $N=K$. Each RF chain is connected to a waveguide, and each waveguide has $L=\left\{8,12,16,20,24,28,32\right\}$ PAs. All the waveguides stretch across a span of $20$ meters along the x-axis (horizontal) direction. Moreover, the waveguides are uniformly spaced along the $y$-axis (vertical) directions with an equal interval of $d_{\mathrm{W}}=S_{\mathrm{y}}/N$ meter. The maximum transmission power is set as $P=\big{\{}10,12,14,16,18,20,22,24\big{\}}$ dBm, and the noise power is $\sigma^{2}=-90$ dBm. The refraction index is given by $n_{\mathrm{eff}}=1.4$. The residual tolerance for the termination of MM-PDD algorithm is set as $\epsilon=10^{-6}$, and the initial penalty factor is $\rho^{(0)}=10^{-4}$. We generate $30000$ samples by Monte-Carlo simulation to train KDL-Transformer, and generate a test dataset of $64$ samples for each simulation setting to compare the performance of the KDL-Transformer and the MM-PDD algorithms. Several baseline algorithms are considered:
  • Massive MIMO (PDD): The conventional massive MIMO array based on hybrid beamforming architecture [28] is adopted at the BS located at the origin point, where PDD algorithm [27] is utilized to jointly optimize the analog beamforming and the transmit beamforming. Aligned with the proposed PASS, the massive MIMO uses a sub-connected structure, where each RF chain is only connected to a subset of $L$ antennas through $L$ phase shifters, thus reducing the hardware costs and complexity.
@@
Fig. 2 demonstrates the convergence behaviors of the developed MM-PDD algorithm in terms of both the achieveable system sum rate and the maximum constraint residual, i.e., $\left\|\mathbf{B}\right\|_{\infty}$. As shown in Fig. 2, by appropriately initializing the penalty factor $\rho$, the developed MM-PDD algorithm can successfully seek the directions to increase the system sum rate and reduce the constraint residuals simultaneously. Specifically, the maximum constraint residuals $\left\|\mathbf{B}\right\|_{\infty}$ rapidly decreases within the first $40$ iterations. As the punishments on constraint residuals become small, the achievable system sum rate will dramatically enhance and finally achieve convergence. At the convergence, the constraint residual eventually approach 00 and becomes stable, which implies that the equality constraints can be satisfied at convergence. When different numbers of PAs $L$ are activated along each waveguide, the developed MM-PDD algorithm can reach convergence within $50$ iterations, which verifies its efficiency. Furthermore, when $L$ increases, the convergence speed would not be decreased, as the punishments on equation violation increase with $L$, which will enforce PAs to quickly determine their locations.
@@
Fig. 3 exhibits the convergence behaviors of the proposed KDL-Transformer algorithm and other learning algorithms, where the eventual performance achieved by MM-PDD algorithm is also displayed for comparisons. The detailed numerical results are further presented in Table I. As shown in Fig. 3, the black-box learning algorithms based on both ResNet and Transformer perform much worse than the MM-PDD algorithm. This is because the coupled beamforming optimization problem has numerous local optimums and saddle points, making it difficult for the unsupervised learning to seek high-quality solutions by solely relying on gradient descent. By contrast, despite the utilized ML models, the proposed KDL method can significantly enhance the performance of learning-based algorithms, which confirms its efficiency to approximate KKT points. Notably, the proposed KDL-Transformer algorithm improves over $30\%$ performance than the MM-PDD algorithm. This demonstrates its potentials to circumvent the inefficiency of iterative alternating block descent required by the optimization algorithms. Furthermore, the proposed KDL-Transformer improves over $40\%$ and over $20\%$ performance then KDL-ResNet and KDL-Transformer-OCA, which confirms that its efficiency in characterizing both inter-PA/inter-user and CSI-beamforming dependencies.
Fig. 4 shows the system performance of different algorithms under various transmitting power limitations $P$. The system sum rates of both the conventional massive MIMO and the proposed PASS framework increase with $P$. Both the MM-PDD algorithm and the KDL-Transformer algorithm significantly improves the system sum rate compared to the conventional massive MIMO architecture, which confirms the efficiency of the proposed PASS architecture in reconfiguring large-scale path loss and spatial multiplexing for MISO communciations. Moreover, the developed KDL-Transformer algorithm achieves the highest system performance, and the achieved performance gain increases with $P$, which demonstrates its potential to enhance the resource utilization efficiency.
@@
Fig. 5 compares the system sum rates of different approaches when different numbers of pinches/antennas are activated on each waveguide/RF chain. Specifically, the system sum rates of both massive MIMO and PASS increase with the number of associated users $K$, which confirms their effectivity in spatial multiplexing. Nevertheless, as the number of activated antennas $L$ increases, the performance of both the massive MIMO and the MM-PDD algorithm deteriorates due to increased mutual interference. By contrast, the performance achieved by KDL-Transformer algorithm may be improved with a larger number of activated pinches. This implies that the optimization-based algorithms may prematurely fall into local optimums, but the pinching beamforming solutions learnt by KDL-Transformer can explore more efficient interference suppression, thus leveraging the the pinching deployment agility when more PAs are activated.
Fig. 6 further demonstrates the performance of different approaches when spatial range $S_{\mathrm{x}}$ varies from $5$ m to $25$ m. It can be observed that the system performance of conventional massive MIMO degrades as the spatial range increases, especially when the transmission power budget is small. In comparison, the performance of the developed MM-PDD algorithm slightly changes as the spatial range increases, since PASS can avoid the long-distance path loss experienced by conventional MIMO. Furthermore, the developed KDL-Transformer consistently outperforms both MM-PDD and conventional massive MIMO, but the performance decreases with the spatial ranges. This may because when the spatial range of user distributions enlarges, the distribution of PAs also becomes more sparse, which will decrease the overall array gains.

### v3

@@ IV-C2 Decoder Design
The decoder dynamically attends to relevant information in the final hidden state $\mathbf{Z}^{\mathrm{E}}$ obtained from the encoder, and predicts the primal/dual variables via a cross-attention mechanism. Let $\mathbf{Z}^{\mathrm{D}}\in\mathbb{R}^{S_{\mathrm{dec}}\times N_{\mathrm{model}}}$ denote the sequence of decoder prediction embeddings, where $S_{\mathrm{dec}}$ is the output sequence length. Specifically, the initial decoder input is constructed as
$\mathbf{Z}^{\mathrm{D}}=\mathbf{Z}_{\mathrm{init}}^{\mathrm{D}}+\mathbf{PE},$ (27)
where $\mathbf{Z}_{\mathrm{init}}^{\mathrm{D}}$ is a learnable parameter matrix and $\mathbf{PE}$ is the sinusoidal positional encoding applied to each decoder element. At each decoding layer, each attention head $i$ of the decoder generates the key matrix for $\mathbf{Z}^{\mathrm{D}}$, while computing query matrix and value matrix using the encoder state $\mathbf{Z}^{\mathrm{E}}$:
$\widetilde{\mathbf{Y}}_{i}^{\mathrm{K}}\!=\!\mathbf{Z}^{\mathrm{D}}\widetilde{\mathbf{W}}_{i}^{\mathrm{K}},\quad\widetilde{\mathbf{Y}}_{i}^{\mathrm{Q}}\!=\!\mathbf{Z}^{\mathrm{E}}\widetilde{\mathbf{W}}_{i}^{\mathrm{Q}},\quad\widetilde{\mathbf{Y}}_{i}^{\mathrm{V}}\!=\!\mathbf{Z}^{\mathrm{E}}\widetilde{\mathbf{W}}_{i}^{\mathrm{V}},$ (28)
where $\widetilde{\mathbf{W}}_{i}^{\mathrm{K}}$, $\widetilde{\mathbf{W}}_{i}^{\mathrm{Q}}$, $\widetilde{\mathbf{W}}_{i}^{\mathrm{V}}$ are trainable decoder parameters. The cross attention score between $\mathbf{Z}^{\mathrm{E}}$ and $\mathbf{Z}^{\mathrm{D}}$ is obtained by:
$\mathrm{head}_{i}\!\!=\!\!CrAttn\!\!\left(\widetilde{\mathbf{Y}}_{i}^{\mathrm{K}},\!\widetilde{\mathbf{Y}}_{i}^{\mathrm{Q}},\!\widetilde{\mathbf{Y}}_{i}^{\mathrm{V}}\right)\!=\!\text{softmax}\!\!\left(\!\tfrac{\widetilde{\mathbf{Y}}_{i}^{\mathrm{Q}}\!\left(\widetilde{\mathbf{Y}}_{i}^{\mathrm{K}}\right)^{\top}}{\sqrt{N_{\mathrm{model}}}}\!\right)\!\widetilde{\mathbf{Y}}_{i}^{\mathrm{V}}\!.$ (29)
@@
$\mathbf{Z}^{\mathrm{D}}\leftarrow\mathbf{Z}^{\mathrm{D}}+\text{MultiHead}\left(\widetilde{\mathbf{Y}}_{i}^{\mathrm{K}},\widetilde{\mathbf{Y}}_{i}^{\mathrm{Q}},\widetilde{\mathbf{Y}}_{i}^{\mathrm{V}}\right).$ (30)
@@
Parameters: Embedding dimension $N_{\text{model}}$, number of heads $I$, numbers of encoder and decoder layers $J_{\mathrm{E}}$ and $J_{\mathrm{D}}$.
1: Compute feature embedding $\mathbf{Z}^{\mathrm{E}}$ by (23).
@@
3:  Compute query $\mathbf{Y}_{i}^{\mathrm{Q}}$, key $\mathbf{Y}_{i}^{\mathrm{K}}$, and value $\mathbf{Y}_{i}^{\mathrm{V}}$ for each attention head $i$ by (24).
4:  Obtain self-attention by (25) and update $\mathbf{Z}^{\mathrm{E}}$ by (26).
@@
6: Initialize decoder output $\mathbf{Z}^{\mathrm{D}}$ by (27).
@@
9:  Update $\mathbf{Z}^{\mathrm{D}}$ by (30).
@@
11: Obtain PA deployment $\mathbf{x}^{\mathrm{end}}\in\mathbb{R}^{N\times 1}$, $\bm{\omega}\in\mathbb{R}^{N\times L}$, dual variable $\bm{\lambda}\in\mathbb{R}^{K\times 1}$, and power allocation $\bm{\mu}\in\mathbb{R}^{K\times 1}$ from $\mathbf{Z}^{\mathrm{D}}=\big[\mathbf{x}^{\mathrm{end}},\bm{\omega},\bm{\lambda},\bm{\mu}\big]\!\in\mathbb{R}^{(N+M+2K)\times 1}$.
12: Project $\mathbf{x}^{\mathrm{end}}$, $\bm{\omega}$ by (31), (33) and obtain $\mathbf{X}$.
@@ IV-C3 Constraint Guarantees
We introduce projection operations in this part to strictly ensure constraints (C1) - (C3). Specfically, the transformer decoder predicts primal and dual variables $\left\{\mathbf{x}^{\mathrm{end}},\bm{\omega},\bm{\mu},\bm{\lambda}\right\}$, and reconstruct feasible solutions of $\mathbf{X}$ and $\mathbf{D}$ through the introduced projections. Vector $\mathbf{x}^{\mathrm{end}}=\big[x_{1}^{\mathrm{end}},x_{2}^{\mathrm{end}},\dots,x_{N}^{\mathrm{end}}\big]^{\top}\in\mathbb{R}^{N\times 1}$ assigns the last PA’s position along each waveguide. Moreover, $\bm{\omega}=\left[\omega_{n,l}\right]\in\mathbb{R}^{N\times L}$ denotes the spacing between any two PAs along the $n$-th waveguide, where $\omega_{n,1}$ is the distance from the first PA to the feed point of waveguide $n$, and $\omega_{n,l}=x_{n,l}-x_{n,l-1}$, $\forall l>1$, indicates the mutual intervals between the remaining PAs.
To satisfy PAs’ range limitation (C2), we first project $x_{n}^{\mathrm{end}}$ into the feasible region $L\Delta_{\min}\leqslant x_{n}^{\mathrm{end}}\leqslant S_{\mathrm{x}}$ using
$x_{n}^{\mathrm{end}}\!\leftarrow\!L\Delta_{\min}\!+\!\mathrm{Sigmoid}\!\left(x_{n}^{\mathrm{end}}\right)\!\left(S_{\mathrm{x}}\!-\!L\Delta_{\min}\right)\!,~\forall n\!\in\!\mathcal{N},$ (31)
where $\mathrm{Sigmoid}(\cdot)$ is the Sigmoid activation function. Further, to simultaneously ensure the minimum spacing constraint (C1), the following projection operation is introduced. Given a threshold $0<\!\epsilon\!\leqslant 1/L$ and a predicted vector $\mathbf{z}=\left[z_{1},z_{2},\dots,z_{L}\right]^{\top}$, $0\!\leqslant\!z_{l}\!\leqslant\!1$, the projection operation
$\mathbf{z}_{\mathrm{proj}}=\mathcal{P}\left(\mathbf{z},\epsilon\right)=\epsilon\mathbf{1}+\frac{1-L\epsilon}{\sum_{l=1}^{L}z_{l}}\mathbf{z},$ (32)
ensures that $z_{\mathrm{proj},l}\!\geqslant\!\epsilon$, $\forall l$, and $\sum\limits_{l=1}z_{\mathrm{proj},l}\!=\!1$. Define $\bm{\omega}_{n}^{T}\!\triangleq\!\big[\omega_{n,1},\allowbreak\omega_{n,2},\dots,\omega_{n,L}\big]\!\in\!\mathbb{R}^{1\times L}$. We enforce the minimum spacing constraints (C1) by applying projection $\mathcal{P}\left(\cdot\right)$ on $\bm{\omega}_{n}$:
$\bm{\omega}_{n}\leftarrow x_{n}^{\mathrm{end}}~\mathcal{P}\left(\bm{\omega}_{n},\tfrac{\Delta_{\min}}{x_{n}^{\mathrm{end}}}\right),~\forall n\in\mathcal{N}.$ (33)
As a result, the $l$-th PA’s position $x_{n,l}$ on waveguide $n$ can be obtained by accumulating the intervals of previous $l$ PAs using $x_{n,l}=\sum_{i=1}^{l}\omega_{n,i},~\forall l\in\mathcal{L},~\forall n\in\mathcal{N}$.
@@
Algorithm 2 summarizes the computation procedure for the developed KDL-Transformer algorithm for predicting solutions. The encoder and the decoder are jointly trained through end-to-end learning to minimize the loss function, i.e., the negative system sum rate. The inference time complexity can be analyzed as follows. The encoder processes user locations with a sequence length of $S_{\mathrm{enc}}\!=\!2K$, while the decoder outputs $\big[\bm{x}_{\mathrm{end}},\bm{\omega},\bm{\lambda},\bm{\mu}\big]$ with a sequence length of $S_{\mathrm{dec}}\!=\!N\!+\!NL\!+\!2K$. Given the embedding width $N_{\mathrm{model}}$ and $J_{E}$ ($J_{D}$) encoder (decoder) layers, the forward propagation of the KDL-Transformer includes: (i) encoder self-attention and feed-forward with complexity $\mathcal{O}\!\big(J_{E}(S_{\mathrm{enc}}^{2}N_{\mathrm{model}}\!+\!S_{\mathrm{enc}}N_{\mathrm{model}}^{2})\big)$; (ii) decoder self-attention, cross-attention, and feed-forward with $\mathcal{O}\!\big(J_{D}(S_{\mathrm{dec}}^{2}N_{\mathrm{model}}\!+\!S_{\mathrm{dec}}N_{\mathrm{model}}^{2}\!+\!S_{\mathrm{dec}}S_{\mathrm{enc}}N_{\mathrm{model}})\big)$; and (iii) KKT-conditioned beamformer reconstruction in (22), requiring $\mathcal{O}(N^{3})$. The overall inference complexity is thus $\mathcal{O}\!\big((J_{E}S_{\mathrm{enc}}^{2}\!+\!J_{D}S_{\mathrm{dec}}^{2}\!+\!J_{D}S_{\mathrm{dec}}S_{\mathrm{enc}})N_{\mathrm{model}}\!+\!(J_{E}S_{\mathrm{enc}}\!+\!J_{D}S_{\mathrm{dec}})N_{\mathrm{model}}^{2}\!+\!N^{3}\big)$.
Remark 4 (Complexity Comparsions):
Considering $L>N=K$ and $S_{\mathrm{dec}}=N+NL+2K$, the inference complexity of KDL-Transformer mainly scales as $\mathcal{O}\big(J_{D}N_{\mathrm{model}}N^{2}L^{2}\big)$. Compared with MM-PDD, whose time complexity is dominated by $\mathcal{O}\big(N_{\mathrm{it}}N^{3}L^{3}\big)$ (typically with $N_{\mathrm{it}}>50$), the KDL-Transformer achieves a much lower cost by fixing $J_{D}$ and $N_{\mathrm{model}}$, such that $J_{D}\ll N_{\mathrm{it}}$ and $N_{\mathrm{model}}\leqslant N^{1.5}L^{1.5}$.
@@
Performance Method $M$-PDD KDL-Transformer KDL-Transformer-OCA KDL-ResNet Black-box Transformer Black-box ResNet Conventional MIMO
@@ V Simulation Result
In this section, we provide numerical results to demonstrate the effectiveness of the proposed PASS design and algorithms, where a BS serves $K=\left\{4,6\right\}$ single-antenna users in downlink MIMO communications. The system operates at $f=30$ GHz frequency band, and the wavelength is $\lambda_{f}=0.01$ meters. Users are randomly distributed in a range of $S_{\mathrm{x}}\times S_{\mathrm{y}}=20\times 10~\text{m}^{2}$ area. Both users and PASS have fixed heights, where $z_{k}^{\mathrm{U}}=0$, $\forall k\in\mathcal{K}$, and $h_{\mathrm{PA}}=2.5$ meter, $\forall n\in\mathcal{N}$. To ensure multiplexing gains, the number of RF chains used by PASS is set as $N=K$. Each RF chain is connected to a waveguide, and each waveguide has $L=\left\{8,12,16,20,24,28,32\right\}$ PAs. All the waveguides stretch across a span of $20$ meters along the x-axis (horizontal) direction [^2] Motion latency of PAs is ignored. To reduce practical timing overheads, a semi-continuous activation [27] can be adopted, which activates pre-installed PAs near the optimized locations via small, fast micro-adjustments.. Moreover, the waveguides are uniformly spaced along the $y$-axis (vertical) directions with an equal interval of $d_{\mathrm{W}}=S_{\mathrm{y}}/N$ meter. The maximum transmission power is set as $P=\big\{10,12,14,16,18,20,22,24\big\}$ dBm, and the noise power is $\sigma^{2}=-90$ dBm. The refraction index is given by $n_{\mathrm{eff}}=1.4$. The residual tolerance for the termination of MM-PDD algorithm is set as $\varepsilon=10^{-6}$. For the initialization of MM-PDD algorithm, PAs are equally spaced around the average user location, $\mathbf{D}^{(0)}$ is initialized by regularized zero-forcing solution, dual variables $\bm{\lambda}^{(0)}=\mathbf{0}$, penalty factor $\rho^{(0)}=10^{-4}$, and other auxiliary variables are randomly initialized. Moreover, $J_{E}=J_{D}=2$, and $N_{\mathrm{model}}=128$. We generate $30000$ samples by Monte-Carlo simulation to train KDL-Transformer [^3] Reproducible code for KDL-Transformer is publicly available at https://github.com/xiaoxiaxusummer/KDL_Transformer_Beamforming.. For each setting of $(P,L,K,S_{x})$, all numerical results are averaged over 64 i.i.d. random test samples. Increasing the size of the test dataset does not noticeably affect the experimental results. Several baseline algorithms are considered:
  • Massive MIMO: The conventional massive MIMO array based on hybrid beamforming architecture [28] is adopted at the BS located at the origin point, where PDD algorithm [29] is utilized to jointly optimize the analog beamforming and the transmit beamforming. Aligned with the proposed PASS, the massive MIMO uses a sub-connected structure, where each RF chain is only connected to a subset of $L$ antennas through $L$ phase shifters, thus reducing the hardware costs and complexity.
@@
Fig. 3 demonstrates the convergence behaviors of the developed MM-PDD algorithm in terms of both the achieveable system sum rate and the maximum constraint residual, i.e., $\left\|\mathbf{B}\right\|_{\infty}$. As shown in Fig. 3, by appropriately initializing the penalty factor $\rho$, the developed MM-PDD algorithm can successfully seek the directions to increase the system sum rate and reduce the constraint residuals simultaneously. Specifically, the maximum constraint residuals $\left\|\mathbf{B}\right\|_{\infty}$ rapidly decreases within the first $40$ iterations. As the punishments on constraint residuals become small, the achievable system sum rate will dramatically enhance and finally achieve convergence. At the convergence, the constraint residual eventually approach 0 and becomes stable, which implies that the equality constraints can be satisfied at convergence. When different numbers of PAs $L$ are activated along each waveguide, the developed MM-PDD algorithm can reach convergence within $50$ iterations, which verifies its efficiency. Furthermore, when $L$ increases, the convergence speed would not be decreased, as the punishments on equation violation increase with $L$, which will enforce PAs to quickly determine their locations.
@@
Fig. 4 exhibits the convergence behaviors of the proposed KDL-Transformer algorithm and other learning algorithms, where the eventual performance achieved by MM-PDD algorithm is also displayed for comparisons. The detailed numerical results are further presented in Table I. As shown in Fig. 4, the black-box learning algorithms based on both ResNet and Transformer perform much worse than the MM-PDD algorithm. This is because the coupled beamforming optimization problem has numerous local optimums and saddle points, making it difficult for the unsupervised learning to seek high-quality solutions by solely relying on gradient descent. By contrast, despite the utilized ML models, the proposed KDL method can significantly enhance the performance of learning-based algorithms, which confirms its efficiency to approximate KKT points. Notably, the proposed KDL-Transformer algorithm improves over $20\%$ performance than the MM-PDD algorithm. This demonstrates its potentials to circumvent the inefficiency of iterative alternating block descent required by the optimization algorithms. Furthermore, the proposed KDL-Transformer improves over $40\%$ and over $18\%$ performance then KDL-ResNet and KDL-Transformer-OCA, which confirms that its efficiency in characterizing both inter-PA/inter-user and CSI-beamforming dependencies.
Fig. 5 shows the system performance of different algorithms under various transmitting power limitations $P$. The system sum rates of both the conventional massive MIMO and the proposed PASS framework increase with $P$. Both the MM-PDD algorithm and the KDL-Transformer algorithm significantly improves the system sum rate compared to the conventional massive MIMO architecture, which confirms the efficiency of the proposed PASS architecture in reconfiguring large-scale path loss and spatial multiplexing for multi-user communciations. Moreover, the developed KDL-Transformer algorithm achieves the highest system performance, and the achieved performance gain increases with $P$, which demonstrates its potential to enhance the resource utilization efficiency.
@@
Fig. 6 compares the system sum rates of different approaches when different numbers of pinches/antennas are activated on each waveguide/RF chain. Specifically, the system sum rates of both massive MIMO and PASS increase with the number of associated users $K$, which confirms their effectivity in spatial multiplexing. Nevertheless, as the number of activated antennas $L$ increases, the performance of both the massive MIMO and the MM-PDD algorithm deteriorates due to increased mutual interference. By contrast, the performance achieved by KDL-Transformer algorithm may be improved with a larger number of activated PAs. This implies that the optimization-based algorithms may prematurely fall into local optimums, but the pinching beamforming solutions learnt by KDL-Transformer can explore more efficient interference suppression, thus leveraging the the pinching deployment agility when more PAs are activated.
Fig. 7 further demonstrates the performance of different approaches as the spatial range $S_{\mathrm{x}}$ varies from $5$ m to $25$ m. For comparison, an ideal fully digital (FD) MIMO system is also included as an additional baseline, where $NL$ antennas are each connected to an RF chain to serve $K$ users, and fully digital beamforming is obtained via the WMMSE algorithm. It can be observed that massive MIMO systems, with the proposed PDD-based hybrid beamforming, can closely approach the average sum rate of FD-MIMO. However, the performance of both massive MIMO and FD-MIMO systems degrades as the spatial range increases, especially under limited transmit power. In contrast, the performance of PASS is slightly affected as the spatial range increases, which demonstrates PASS’s unique capability to adjust large-scale path loss. Furthermore, although the KDL-Transformer outperforms MM-PDD on average, it is less robust across varying spatial ranges. This is likely because a larger spatial range leads to a sparser PA distribution and reduced array gain using KDL-Transformer.